OpenStudy (anonymous):

Rachel mows lawns. Her lawn mower cuts a path of 2 feet wide, and moves forward at an average speed of 180 feet per minute. Rachel is mowing a rectangular section of lawn that measures 120 feet long and 80 feet wide. Make a reasonable estimate for how long it will take her to mow this section on lawn. Rachel started mowing the lawn at 3:00 pm. After 10 minutes of mowing she increased her speed to 200 feet per minute. Approximately what time will Rachel finish mowing the lawn? Please show your work or explain your answer.

OpenStudy (anonymous):

120*80 = 9600 which is the area. every second she moves 180/60 = 3 feet forward. so every second she mows a block of grass that is 3x2 = 6 feet in area. 9600/6 = 1600 seconds = 26.7 minutes. enjoy

OpenStudy (anonymous):

How do I finish the second part of the question?